OpenStudy (anonymous):

Write the product as a trinomial. (m – 5)(m + 2) A. m2– 3m + 10 B. m2 – 10 C. m2 + 3m – 10 D. m2 – 3m – 10

OpenStudy (primeralph):

FOIL?

OpenStudy (anonymous):

Yeah

OpenStudy (primeralph):

So FOIL.

OpenStudy (jadeishere):

m^2 + 2m - 5m - 10 m^2 - 3m - 10

OpenStudy (anonymous):

I know it's not b, But i have a hard time deciding on what to add and subtract..

OpenStudy (anonymous):

(m – 5)(m + 2) =(mxm)+(mx2)+(-5xm)+(-5x2) =m^2-3m-10 do you understand what i did?